List of Photos

1 placement left. abc abc abc abc abc abc abc abc abc abc abc abc abc abc abc abc

abc abc abc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2 placement right. abc abc abc abc abc abc abc abc abc abc abc abc abc abc abc abc

abc abc abc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3 placement inside. abc abc abc abc abc abc abc abc abc abc abc abc abc abc abc

abc abc abc abc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 4 placement outside. abc abc abc abc abc abc abc abc abc abc abc abc abc abc abc

abc abc abc abc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 5 placement outside. abc abc abc abc abc abc abc . . . . . . . . . . . . . . . . . . . 2 6 placement outside. t caption. abc abc abc abc abc abc abc . . . . . . . . . . . . . 3 7 placement outside. c caption. abc abc abc abc abc abc abc . . . . . . . . . . . . . 3 8 placement outside. short c caption. . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 9 placement outside. b caption. abc abc abc abc abc abc abc . . . . . . . . . . . . . 3 10 placement inside. t caption. u. abc abc abc abc abc abc abc . . . . . . . . . . . . 3 11 placement inside. c caption. u. abc abc abc abc abc abc abc . . . . . . . . . . . . 3 12 inside. short c caption. u. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 13 placement inside. b caption. u. abc abc abc abc abc abc abc . . . . . . . . . . . . 3 14 photo l. t caption. taker s. abc abc abc abc abc abc abc . . . . . . . . . . . . . . 4 15 photo l. c caption. taker s. abc abc abc abc abc abc abc . . . . . . . . . . . . . . 4 16 photo l. c caption. taker s. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 17 photo l. b caption. taker s. abc abc abc abc abc abc abc . . . . . . . . . . . . . . 4 18 photo r. t caption. taker s. abc abc abc abc abc abc abc . . . . . . . . . . . . . . 4 19 photo r. c caption. taker s. abc abc abc abc abc abc abc . . . . . . . . . . . . . . 4 20 photo r. c caption. taker s. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 21 photo r. b caption. taker s. abc abc abc abc abc abc abc . . . . . . . . . . . . . . 4 22 defaults. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 23 lop caption . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 24 Photo environment. i. caption for this. . . . . . . . . . . . . . . . . . . . . . . . . . 5 25 Photo env lop cap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 26 defaults again. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

Mg_{grapher}

Photo 1: placement left. abc abc abc abc abc abc abc abc abc abc abc abc abc abc abc abc abc abc abc

Mg_{grapher}

Photo 2: placement right. abc abc abc abc abc abc abc abc abc abc abc abc abc abc abc abc abc abc abc

Mg

photographer

Photo 3: placement inside. abc abc abc abc abc abc abc abc abc abc abc abc abc abc abc abc abc abc abc

Mg

photographer

Photo 4: placement outside. abc abc abc abc abc abc abc abc abc abc abc abc abc abc abc abc abc abc abc

Mg

photographer

Photo 5: placement outside. abc abc abc abc abc abc abc Photo 6: placement outside. t caption. abc abc abc abc abc abc abc Mg

photographer

Photo 7: placement outside. c caption. abc abc abc abc abc abc abc Mg

photographer

Photo 8: placement outside. short c caption. Mg

photographer

Photo 9: placement outside. b caption. abc abc abc abc abc abc abc Mg

photographer

Mg Photo 10: placement inside. t caption. u. abc abc abc abc abc abc abc

photographer

Mg Photo 11: placement inside. c caption. u. abc abc abc abc abc abc abc

photographer

Mg  Photo 12: inside. short c caption. u.

photographer

Mg

Photo 13: placement inside. b caption. u. abc abc abc abc abc abc abc Mg Photo 14: photo l. t caption. taker s. abc abc abc abc abc abc abc

photographer

photographer

Mg Photo 15: photo l. c caption. taker s. abc abc abc abc abc abc abc

photographer

Mg  Photo 16: photo l. c caption. taker s.

photographer

Mg

Photo 17: photo l. b caption. taker s. abc abc abc abc abc abc abc

photographer

photographer^{Photo 18: photo r. t caption. taker s. abc abc abc abc abc abc abc Mg}

photographer^{Photo 19: photo r. c caption. taker s. abc abc abc abc abc abc abc Mg}

photographer^{Photo 20: photo r. c caption. taker s.  Mg}

photographer Photo 21: photo r. b caption. taker s. abc abc abc abc abc abc abc Mg

Photo 22: defaults.  Mg

photographer

Photo 23: lop cap, defaults.  Mg

photographer

Mg  Photo 24: Photo environment. i. caption for this.

phogr

Mg  Photo 25: Photo environment. i. caption.

phogr

Photo 26: defaults again.  Mg

photographer